you have 8 shirts and plan to wear a different one each day from Monday through Friday how many possible arrangements of shirts are possible for those 5 days?
Answer by Herdaysoji(25)

8C5 .'.combination which is nCr=(n!)/((r!)(n-r)!)
8!/(5!)(8-5)!=8!/(5!)(3!)=(8*7*6)/(3*2*1)
336/6=56
.'. it is 56 types of arrangements